# Moving Aggregate Modified Autoregressive Copula-Based Time Series Models (MAGMAR-Copulas)

### v2

@@ 3.2 Relation to other models
Now we derive the relation of the proposed MAGMAR-Copula model to the classical copula-based time series model and the classical linear ARMA model. The results are given for the MAGMAR(1,1)11(1,1)( 1 , 1 ) case but can be extended to the MAGMAR(1,1)11(1,1)( 1 , 1 ) case. Furthermore we define the MAG-copula time series and show that it is stationary. The first two proposition show that the classical linear ARMA model as well as the classical copula-based time series model is nested by the MAGMAR-Copula model.
@@
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. If the AR-Copula, $C_{\phi}$, and the MAG-Copula, $C_{\theta}$, are chosen as the normal copula, $C_{\phi}=C^{\text{normal}}_{\phi}$, $C_{\theta}=C^{\text{normal}}_{\theta}$ with $|\phi|<1$ and $|\theta|<1$, the quantile transformed time series $\{\varphi^{-1}(U_{t})\}_{t\in\mathbb{Z}}$ is the classical linear ARMA(1,1)11(1,1)( 1 , 1 ) time series model with AR-parameter $\phi$, MA-parameter $\frac{\beta}{(1-\beta^{2})^{\frac{1}{2}}}$ and independent normally distributed innovations with variance $(1-\phi^{2})(1-\theta^{2})$.
@@
See Appendix A
Proposition 2.
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. If the MAG-copula is the independence copula, $C_{\theta}=\Pi$, then the classical copula-based time series model (see Sect. 2.1) is recovered.
Proof.
The $h$-function associated to the independence copula is $h_{\Pi}(u_{1},u_{2})=\frac{\partial(u_{1}u_{2})}{\partial u_{2}}=u_{1}$. The corresponding inverse $h$-function is $h_{\Pi}^{-1}(u_{1},u_{2})=u_{1}$. It follows that $h_{\theta}^{-1}(w_{t},w_{t-1})=w_{t}$, so the updating equation becomes $U_{t}=h_{\phi}^{-1}(w_{t},U_{t-1}),w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$, which is the updating equation of the classical copula-based time series model (see Eq. 10). ∎
This result is expected, since the proposed MAGMAR-Copula model is a generalization of the classical copula-based time series model. This result also justifies, describing the copula-based time series model as MAGMAR$(1,0)$-Copula model in further discussions. Next we define the MAG-Copula model.
@@
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. If the AR-copula is set to be the independence copula, $C_{\phi}=\Pi$, the resulting model is the MAG(1)-Copula model.
We note again that the MAG$(1)$-Copula model coincides with the $1$-dependent model from Chapter 3.14 of Joe (2014). It is easy to see that that a MAG(1)-Copula time series, $\{V_{t}\}_{t\in\mathbb{Z}}$ follows the updating equation
@@
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAG$(1)$ time series. Then $\{U_{t}\}_{t\in\mathbb{Z}}$ is strictly stationary and $U_{t}\sim U(0,1),\ \forall t\in\mathbb{Z}$.
@@
See Appendix C
The models distributional properties are important for further work. They play an important role in establishing the stationarity of the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model with the normal AR-Copula, see Sect. 3.3.1. The next proposition treats the trivial case, where both, the AR and the MAG copula are chosen as the independence copula.
@@
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. If the AR and the MAG-copula are set to be the independence copula, $C_{\theta}=C_{\phi}=\Pi$ the time series $\{U_{t}\}_{t\in\mathbb{Z}}$ is an $iid$ standard uniform time series.
Proof.
Again we use that the inverse $h$-function associated to the independence copula is $h_{\Pi}^{-1}(u_{1},u_{2})=u_{1}$. From this it follows that the model updating equation is $U_{t}=w_{t}$, where $w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$. ∎
3.3 Distributional Properties of the model
In this section, the distributional properties of random variables $\{U_{i}\}_{i\in\mathbb{Z}}$, following the MAGMAR-Copula model, are derived. In order to keep the findings comprehensible, the ideas are developed for the MAGMAR(1,1)11(1,1)( 1 , 1 ) case. They can be generalized to the MAGMAR($p,q$) case.
First, we show that process is stationary for the case where the AR-copula is the normal copula and the parameter fulfills $|\phi|<1$. Second, we derive the conditional distribution of $U_{t}$ given past obervations. Third, using the law of total probability and conditional independence we derive the joint distribution of $\{U_{i}\}_{i\in\mathbb{Z}}$.
3.3.1 Stationarity
We show that the process is stationary for the case where the AR-copula is the normal copula. In order to do so, we consider the MAG representation of the MAGMAR$(1,0)$-normal-Copula time series and show that it does converge almost surely for all $t$. Together with the stationarity of the MAG$(1)$-Copula model, the stationarity of the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model will follow.
Proposition 5.
Let $\{U_{t}\}$ be a MAGMAR$(1,0)$-Copula time series where the AR-Copula is the normal copula, $C_{\phi}=C_{\phi}^{\text{normal}}$, with parameter $|\phi|<1$. Then the MAG()($\infty$)( ∞ ) representation,
@@ 3.3.1 Stationarity
The inverse $h$-function associated to a normal copula with parameter $\phi$ is given as $h_{\phi}^{-1}(u_{1},u_{2})=\varphi(\varphi^{-1}(u_{1})(1-\phi^{2})^{\frac{1}{2}}+\phi\varphi^{-1}(u_{2}))$ (Smith et al., 2010). Then $A_{t,n}$ can be written as
$A_{t,n}=\varphi\left((1-\phi^{2})^{\frac{1}{2}}\sum_{i=0}^{n}\phi^{i}\varphi^{-1}(w_{t-i})\right).$
Since $w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U[0,1]$ and $\phi^{i}$ is decreasing geometrically in $i$, the variance of the sum $\sum_{i=0}^{n}\phi^{i}\varphi^{-1}(w_{t-i})$ is bounded and it converges almost surely (Billingsley, 1995). The limiting random variable is distributed as $N(0,\frac{1}{1-\phi^{2}})$. By the continuous mapping theorem and the continuity of $\varphi$ as well as the continuity of the multiplication with $(1-\phi^{2})^{\frac{1}{2}}$ it follows that $A_{t,n}$ also converges almost surely. Since $(1-\phi^{2})^{\frac{1}{2}}\sum_{i=0}^{n}\phi^{i}\varphi^{-1}(w_{t-i})\sim N(0,1)$, the limiting random variable is standard uniformly distributed. ∎
Notice that the same result holds when instead of an $iid$ sequence $w_{t}$, a dependent, but stationary, sequence is plugged in. This is in line with Theorem 36.4 from Billingsley (1995). Furthermore, from Lemma 5 it follows that the MAGMAR$(1,0)$-Copula time series is strictly stationary. We want to note, that the MAG($\infty$)-representation of the MAGMAR$(1,0)$-Copula time series is analogous to the MA($\infty$) representation in the classical linear ARMA model (see Hamilton, 2020) and is equivalent to the representation derived in McNeil and Bladt (2022). In principle the MA($\infty$) representation of the classical linear ARMA model can also be written as a infinite composition. However, there the composed functions are linear and, hence, also their composition is linear. We noted earlier that the $a.s.$ convergence in proposition 5 also holds if the innovation time series $\{w_{t}\}_{t\in\mathbb{Z}}$ is not $iid$ but merely stationary. In the next proposition we exploit this fact to show that the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series with the AR-Copula being the normal copula is stationay.
@@
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series where the AR-Copula is the normal copula, $C_{\phi}=C_{\phi}^{\text{normal}}$, with parameter $|\phi|<1$. Then $\{U_{t}\}_{t\in\mathbb{Z}}$ is strictly stationary.
@@
Similar to the reasoning in the proof of Lemma 5, the MAG()($\infty$)( ∞ ) representation of $\{U_{t}\}$ can be written as
$A_{t,n}=\varphi\left((1-\phi^{2})^{\frac{1}{2}}\sum_{i=0}^{n}\phi^{i}\varphi^{-1}(h_{\theta}^{-1}(w_{t-i},w_{t-i-1}))\right).$
Now notice that $\{V_{t},\ V_{t}:=h_{\theta}^{-1}(w_{t},w_{t-1})\}_{t\in\mathbb{Z}}$ is a MAG$(1)$ time series. From proposition 3 it follows that $\{V_{t}\}_{t\in\mathbb{Z}}$ is strictly stationary, so $\{\varphi^{-1}(V_{t})\}_{t\in\mathbb{Z}}$ is also strictly stationary. Since by proposition 5 converges almost surely and thereby constitutes a measurable mapping, by Theorem 36.4 of Billingsley (1995) it follows that the process is strictly stationary. ∎
Note that the stationary distribution of $\{U_{t}\}_{t\in\mathbb{Z}}$ does not necessarily have to be uniform. This may be problematic for forecasting and general modeling. A possible solution would be to apply the PIT to the random variables on the pseudo-observation scale before applying the quantile transformation. We would like to explore this in future research. Next, we derive the conditional distribution of $U_{t}$ given all past observations.
@@ 3.3.2 Conditional distribution
The conditional distribution of $U_{t}|\mathcal{F}_{t-1}$, where $\mathcal{F}_{s}:=\sigma(\{X_{r}\}_{r\leq s})$, can be derived by using the density transformation theorem (see e.g. Devore et al., 2012, Sect. 4.7). We can interpret the updating equation of the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model conditional on $(U_{t-i})_{i=1}^{\infty}:=(U_{t},U_{t-1},\ldots)$ as follows: The random variable $w_{t}\sim U(0,1)$ is transformed by a function $g:(0,1)\rightarrow(0,1)$ to the random variable $U_{t}$, so $U_{t}|\mathcal{F}_{t-1}=g(w_{t})|\mathcal{F}_{t-1}$. The other random variables, $(U_{t-i})_{i=1}^{\infty}$, are regarded as deterministic due to conditioning. In the function $g$, $(U_{t-i})_{i=1}^{\infty}$ are regarded as constants. We give an explicit form for $g$ in the proof of the following proposition.
@@
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. Furthermore, let the AR-Copula, $C_{\phi}$, and the MAG-Copula, $C_{\theta}$, be differentiable. Then the conditional density of $U_{t}$ given $\mathcal{F}_{t-1}$ is given as
@@
where $c_{\theta}$ is the copula density associated to the MAG-Copula and $c_{\phi}$ is the density associated to the AR-Copula. The expression $w_{t-1}((U_{t-i})_{i=1}^{\infty})$ is to be understood as a function of conditioning variables $(U_{t-i})_{i=1}^{\infty}$, which are regarded as constant in the conditional density.
@@
Let $g:(0,1)\rightarrow(0,1)$ be given as
$g(x)=$ $h_{\phi}^{-1}\left(h_{\theta}^{-1}(x,w_{t-1}((U_{t-i})_{i=1}^{\infty}),U_{t-1}\right).$
where $w_{t-1}((U_{t-i})_{i=1}^{\infty})$ is again a function of the conditioning variables $(U_{t-i})_{i=1}^{\infty}$. The explicit form of $w_{t}((U_{t-i})_{i=1}^{\infty})$ is given by the AR($\infty$) representation, Eq. 14. It holds that $U_{t}|\mathcal{F}_{t-1}=g(w_{t})|\mathcal{F}_{t-1}$. Application of the density transformation theorem (Devore et al., 2012, Sect. 4.7) yields the result. (The calculation of the Jacobian determinant is given in Appendix B) ∎
It is evident from the conditional density, Eq. 17, that $U_{t}|\mathcal{F}_{t-1}=U_{t}|(U_{t-i})_{i=1}^{\infty}=U_{t}|(U_{t-1},w_{t-1})$. This means that conditioning on all past random variables $(U_{t-i})_{i=1}^{\infty}$ is equivalent to conditioning on $(U_{t-1},w_{t-1})$. This is the same property as in the classical linear ARMA model. Note that for the special case where $c_{\theta}$ is the independence copula density, $c_{\theta}=1_{(0,1)^{2}}$, the conditional density of the copula-based time series model with autoregressive order of one is recovered. In the generalized MAGMAR($p,q$) case, the conditional density can be derived in a similar fashion. The random variable $w_{t}$ is transformed to the random variable $U_{t}$ by an appropriate transformation function $g$. The resulting density is a rather complicated expression which is given in Appendix E. Furthermore we note that multivariate conditional densities such as $f_{U_{t},U_{t-1}|\mathcal{F}_{t-2}}$ can be derived in a similar way. The derivation of joint conditional densities for the MAGMAR(1,1)11(1,1)( 1 , 1 ) model can be found in Appendix F.
3.3.3 Joint distribution/Likelihood
@@
[…]

### v3

@@ 3.2 Relation to other models
Now we derive the relation of the proposed MAGMAR-Copula model to the classical copula-based time series model and the classical linear ARMA model. The results are given for the MAGMAR(1,1)(1,1) case but can be naturally extended to the MAGMAR(p,q)(p,q) case. Furthermore we define the MAG-copula time series and show that it is stationary.
@@
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)(1,1)-Copula time series with AR-Copula $C_{\phi}$ and the MAG-Copula, $C_{\theta}$. The following holds.
  • i)
    If $C_{\phi}$ and $C_{\theta}$ are chosen as the normal copulas, the quantile transformed time series $\{\Phi^{-1}(U_{t})\}_{t\in\mathbb{Z}}$ is the classical Gaussian and linear ARMA(1,1)(1,1) model with AR-parameter $\phi$, MA-parameter $\frac{\theta}{(1-\theta^{2})^{\frac{1}{2}}}$ and independent normally distributed innovations with variance $(1-\phi^{2})(1-\theta^{2})$.
  • ii)
    If the MAG-copula is the independence copula, $C_{\theta}=\Pi$, then the classical copula-based time series model (see Sect. 2.2) is recovered.
  • iii)
    If the AR-copula is the independence copula, then the MAG(1)(1)-Copula time series is recovered.
  • iv)
    If both the AR and MAG-copula are chosen as the independence copula, the time series $\{U_{t}\}_{t\in\mathbb{Z}}$ is an iidiid standard uniform time series.
@@
See Appendix B.1. ∎
The first two statements show that the classical linear ARMA model as well as the classical copula-based time series model is nested by the MAGMAR-Copula model. Part ii) also justifies describing the copula-based time series model as MAGMAR(1,0)(1,0)-Copula model in further discussions. Next we define the MAG-Copula model.
@@
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)(1,1)-Copula time series. If the AR-copula is set to be the independence copula, $C_{\phi}=\Pi$, the resulting model is the MAG(1)-Copula model.
We note again that the MAG(1)(1)-Copula model coincides with the $1$-dependent model from Chapter 3.14 of Joe (2014). It is easy to see that that a MAG(1)-Copula time series, $\{V_{t}\}_{t\in\mathbb{Z}}$ follows the updating equation
@@
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAG(1)(1) time series. Then $\{U_{t}\}_{t\in\mathbb{Z}}$ is strictly stationary, ergodic and $U_{t}\sim U(0,1),\ \forall t\in\mathbb{Z}$.
@@
See B.2
The models distributional properties are important for further work. They play an important role in establishing the stationarity of the MAGMAR(1,1)(1,1)-Copula model with the normal AR-Copula, see Sect. 3.3.1.
3.3 Distributional Properties of the model
In this section, the distributional properties of random variables $\{U_{i}\}_{i\in\mathbb{Z}}$, following the MAGMAR-Copula model, are derived. In order to keep the findings comprehensible, the ideas are developed for the MAGMAR(1,1)(1,1) case. They can be generalized to the MAGMAR($p,q$) case. First, we show that the process is stationary for the case where the AR-copula is the normal copula and the parameter fulfills $|\phi|<1$. Second, we derive the conditional distribution of $U_{t}$ given past observations. Third, using the law of total probability and conditional independence we derive the joint distribution of $\{U_{i}\}_{i\in\mathbb{Z}}$.
3.3.1 Stationarity
We first show that the process is stationary for the case where the AR-copula is the normal copula. Then we generalize the result for the case, where the AR-copula fulfills a contraction-on-average property. In order to establish stationarity for the normal AR-copula case, we consider the MAG representation of the MAGMAR(1,0)(1,0)-normal-Copula time series and show that it does converge almost surely for all $t$. Together with the stationarity of the MAG(1)(1)-Copula model, the stationarity of the MAGMAR(1,1)(1,1)-Copula model will follow. The result is a special case of the more general result on MAGMAR(1,0)(1,0) models, where the copula on average induces a contraction in proposition 5. We still propose the result for the normal AR-copula for illustrational purpose
@@ 3.3.1 Stationarity
Let $\{U_{t}\}$ be a MAGMAR(1,0)(1,0)-Copula time series where the AR-Copula is the normal copula, $C_{\phi}=C_{\phi}^{\text{normal}}$, with parameter $|\phi|<1$. Then the MAG()($\infty$) representation,
@@
See B.3
Notice that the almost sure convergence also holds when instead of an iidiid sequence $w_{t}$, a dependent, but stationary, sequence is plugged in. This is in line with Theorem 36.4 from Billingsley (1995). Furthermore, from Lemma 3 it follows that the MAGMAR(1,0)(1,0)-Copula time series is strictly stationary. We want to note, that the MAG($\infty$)-representation of the MAGMAR(1,0)(1,0)-Copula time series is analogous to the MA($\infty$) representation in the classical linear ARMA model (see Hamilton, 2020) and is equivalent to the representation derived in McNeil and Bladt (2022). In principle the MA($\infty$) representation of the classical linear ARMA model can also be written as a infinite composition. However, there the composed functions are linear and, hence, also their composition is linear. We noted earlier that the $a.s.$ convergence in proposition 3 also holds if the innovation time series $\{w_{t}\}_{t\in\mathbb{Z}}$ is not iidiid but merely stationary. In the next proposition we exploit this fact to show that the MAGMAR(1,1)(1,1)-Copula time series with the AR-Copula being the normal copula is stationay.
@@
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)(1,1)-Copula time series where the AR-Copula is the normal copula, $C_{\phi}=C_{\phi}^{\text{normal}}$, with parameter $|\phi|<1$. Then $\{U_{t}\}_{t\in\mathbb{Z}}$ is strictly stationary.
@@
See B.4
Now we generalize the result from proposition 4 for general bivariate copulas that fulfill a contraction-on-average assumption.
Proposition 5 (AR-Copula contraction).
[…]
Proof.
See B.5
The conditions have to be checked for each AR-copula individually. In particular, a suitable mapping $\kappa$ has to be found. In A.5 we check the conditions for the normal, $t$ and Clayton copula. The generalization of the result for $q>1$ is straight-forward, since the resulting MAG(q)(q) process is still stationary and ergodic. The generalization for $p>1$ is not straight-forward. The typical approach would be to extend the updating equation to a $p$-dimensional updating equation and to show that the corresponding mapping induces a contraction. This approach is not straight-forward for copula-based models because the inverse Rosenblatt transformation is hard to tract and it is not clear how to establish a contraction. This question will be the subject of future research. Note that the proposition now enables us to consider the adjustment transformation $\Psi$ from Def. 1 as the unconditional stationary distribution function of the process, since we now proved such a distribution exists under suitable conditions. Now we proceed to derive the conditional distributions of the model.
@@ 3.3.2 Conditional distribution
The conditional distribution of $U_{t}|\mathcal{F}_{t-1}$, where $\mathcal{F}_{s}:=\sigma(\{X_{r}\}_{r\leq s})$, can be derived by using the density transformation theorem (see e.g. Devore et al., 2012, Sect. 4.7). We can interpret the updating equation of the MAGMAR(1,1)(1,1)-Copula model conditional on $\mathbf{U}_{t-1}:=(U_{t-i})_{i=1}^{\infty}=(U_{t},U_{t-1},\ldots)$ as follows: The random variable $w_{t}\sim U(0,1)$ is transformed by a function $g_{\mathbf{U}_{t-1}}:(0,1)\rightarrow(0,1)$ to the random variable $U_{t}$, i.e. $U_{t}|\mathcal{F}_{t-1}=g_{\mathbf{U}_{t-1}}(w_{t})|\mathcal{F}_{t-1}$. The other random variables, $\mathbf{U}_{t-1}=(U_{t-i})_{i=1}^{\infty}$ , are regarded as deterministic due to conditioning. We give an explicit form for $g_{\mathbf{U}_{t-1}}$ in the proof of the following proposition.
@@
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)(1,1)-Copula time series. Furthermore, let the AR-Copula, $C_{\phi}$, and the MAG-Copula, $C_{\theta}$, be differentiable. Then the conditional density of $U_{t}|\mathcal{F}_{t-1}$ is given as
@@
where $c_{\theta}$ is the copula density associated to the MAG-Copula and $c_{\phi}$ is the density associated to the AR-Copula. The expression $w_{t-1}((U_{t-i})_{i=1}^{\infty})$ is to be understood as a function of conditioning variables $(U_{t-i})_{i=1}^{\infty}$, which are regarded as constant in the conditional density. For the adjusted time series, $\{\tilde{U}_{t}\}_{t\in\mathbb{Z}}=\{\Psi(U_{t})\}_{t\in\mathbb{Z}}$, the conditional density is given as
$f_{\tilde{U}_{t}|\mathcal{F}_{t-1}}(x)=$ $c_{\theta}\left(h_{\phi}(\Psi^{-1}(x),\Psi^{-1}(\tilde{U}_{t-1})),w_{t-1}(\{\tilde{U}_{t-i}\}_{i=1}^{\infty})\right)c_{\phi}(\Psi^{-1}(x),\Psi^{-1}(\tilde{U}_{t-1}))$
$\times\frac{1}{\psi(\Psi^{-1}(x))},$
where $\psi$ is the density corresponding to the distribution function $\Psi$.
@@
See B.6
It is evident from the conditional density, Eq. 16, that $U_{t}|\mathcal{F}_{t-1}=U_{t}|(U_{t-i})_{i=1}^{\infty}=U_{t}|(U_{t-1},w_{t-1})$. This means that conditioning on all past random variables $(U_{t-i})_{i=1}^{\infty}$ is equivalent to conditioning on $(U_{t-1},w_{t-1})$. This is the same property as in the classical linear ARMA model. Note that for the special case where $c_{\theta}$ is the independence copula density, $c_{\theta}=1_{(0,1)^{2}}$, the conditional density of the copula-based time series model with autoregressive order of one is recovered. In the MAGMAR(p,q)(p,q) case, the conditional density can be derived in a similar fashion. The random variable $w_{t}$ is transformed to the random variable $U_{t}$ by an appropriate transformation function $g_{\mathbf{U}_{t-1}}$. The resulting density is a rather complicated expression which is given in A.3. Furthermore we note that multivariate conditional densities such as $f_{U_{t},U_{t-1}|\mathcal{F}_{t-2}}$ can be derived in a similar way. The derivation of joint conditional densities for the MAGMAR(1,1)(1,1) model can be found in A.4.
3.3.3 Joint distribution and Likelihood
@@
For the MAGMAR(1,1)-Copula model, the conditional distribution from Eq. 16 can be plugged in for $f_{U_{t-i}|\mathcal{F}_{t-i-1}}$. For the MAGMAR(p,q)(p,q)-model the conditional densities from A.3, Eq. 20 need to be used. In spite of the factorization, the infinite partial dependence is carried by the $\{w_{i}\}_{i=1}^{T}$. For maximum likelihood estimation purposes, the initial density, $f_{U_{1}}$, may be dismissed, hence resulting in pseudo-MLE. It is easy to verify (via integration by substitution) that $\int f_{U_{t},\ldots,U_{1}}(u_{t},\ldots,u_{1})du_{t}=f_{U_{t-1},\ldots,U_{1}}$, which is to be expected of a joint density. Using the joint density derived in this section, maximum likelihood estimation can be performed. This is discussed in the next section.